What are the values for x that satisfy the equation (x+a) (x+b) =0? F. -a and -b G. -a and b H. -ab J. a and -b K. a and b
Answer by jim_thompson5910(35256)

(x+a)(x+b) =0

x+a=0 or x+b=0

x+a-a=0-a or x+b-b=0-b

x = -a or x = -b

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------